OpenStudy (anonymous):

how can I find the point of intersection of two lines 3x-5y+22=0 and 10x+6y-40=0 ( which are perpendicular)

OpenStudy (precal):

Use substitution. 3x-5y=22 is y=(22/-5) +(22/3)X sub that y equation into 10x+6y-40=0 10x +69((22/-5)+(22/3)X)-40=0 solve for y.